OpenStudy (anonymous):

if A and B as shown below are two mutally exclusive events, what P(A or B)? P(A)=1/8,P(B)=2/5

OpenStudy (danjs):

A or B - either one happens

OpenStudy (anonymous):

the possible choices are 21/40, 3/20, 3/13, and 1/20

OpenStudy (danjs):

So this is the defenition --

OpenStudy (anonymous):

what do you mean? @DanJS

OpenStudy (danjs):

Rule of Addition The probability that Event A or Event B occurs is equal to the probability that Event A occurs plus the probability that Event B occurs minus the probability that both Events A and B occur. P(A ∪ B) = P(A) + P(B) - P(A ∩ B)) Note: Invoking the fact that P(A ∩ B) = P( A )P( B | A ), the Addition Rule can also be expressed as P(A ∪ B) = P(A) + P(B) - P(A)P( B | A )